OpenStudy (anonymous):

AP Calculus: Find the number of units x that produce a maximum revenue R. R=600x^2 - .02x^3

OpenStudy (anonymous):

x=20000 to maximize set the first derivative equal zero.. then you can solve it easily

OpenStudy (anonymous):

1200x-.06x^2=0

OpenStudy (anonymous):

so how did u get 20000

OpenStudy (anonymous):

x(1200-0.06x)=0 then u set both part equal zero so u have the minimum and maximum for x... x=0 and 1200-0.06x=0 ... 1200/0.06=x

OpenStudy (anonymous):

ty can u help me with another problem?

OpenStudy (anonymous):

yes i can try

OpenStudy (anonymous):

k ima open it in a new question

OpenStudy (anonymous):

ok